OpenStudy (anonymous):

Line CD contains points A (4, −7) and B (4, 8). The slope of line CD is zero undefined positive negative

OpenStudy (imstuck):

Do you know slope formula?

OpenStudy (imstuck):

It is:\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]which is rise over run.

OpenStudy (imstuck):

You get these x and y values from your two points (coordinates). y2 is 8, y1 is -7, x2 is 4, and x1 is 4. So the equation is set up like this...

OpenStudy (imstuck):

\[\frac{ 8-4 }{ 4-4 }=\frac{ 4 }{ 0 }\]And of course you know that having a 0 in the denominator of a fraction is mathematically illegal, so the slope is undefined. That means that it is a straight up and down line.